OpenStudy (anonymous):

For what values of a and b is the line 2x+y=b tangent to the parabola y = ax^2 when x = 2?

myininaya (myininaya):

Have you found y'?

OpenStudy (anonymous):

of which ONE

myininaya (myininaya):

Find y' and put 2x+y=b into this form: y=mx+c to find m Then set y'=m

OpenStudy (anonymous):

like y' = 2ax? that one? i start there?

myininaya (myininaya):

ok now set y'=the slope of the tangent line

OpenStudy (anonymous):

uh you mean like 2x + y = b ==> y = -2x+b .. ?

myininaya (myininaya):

and also we want to know what a is when x=2 so we have 2a(2)=the slope of the tangent line

OpenStudy (anonymous):

oy

myininaya (myininaya):

y=-2x+b so the slope of tangent line is -2 right?

myininaya (myininaya):

\[2a(2)=-2\] solve for a! :)

OpenStudy (anonymous):

ah

myininaya (myininaya):

\[a=-\frac{1}{2}\] we also need to find b though!

myininaya (myininaya):

when x=2, what is f(2)?

myininaya (myininaya):

i will let you tell me

myininaya (myininaya):

f(x)=ax^2 we found a what is f(2)?

myininaya (myininaya):

The reason we need this is so we can know a point on the tangent line so we can solve for b

myininaya (myininaya):

2x+y=b 2(2)+f(2)=b 4+f(2)=b so once you find f(2) you will be able to find b

myininaya (myininaya):

Right We could say that with out the bad language part though please

myininaya (myininaya):

without*